### v3

# A Categorical Generalization of Counterpoint

Abstract.
We extend Mazzola’s counterpoint model using category theory, generalizing from the category $\mathbf{Set}$ to other topoi with suitable properties. This generalization suggests that counterpoint’s essential structure depends on specific categorical conditions rather than classical set-theoretic reasoning.
A key contribution is identifying the minimal requirements for counterpoint theory: the topos satisfying some version of Zorn’s Lemma (ZL) and being two-valued with split supports (NS).
Within this framework, we introduce (weak) quasidichotomies alongside the classical notion of dichotomy. These structures capture varying degrees of oppositional structure between consonance and dissonance, with weak quasidichotomies preserving the non-Boolean flexibility essential to musical practice while quasidichotomies represent maximal opposition short of complete partition.
We prove a generalized counterpoint theorem: sequences of admitted successors exist in any topos satisfying our conditions. The framework naturally accommodates counterpoint with sets instead of pure pitches, relaxing the “yes/no” character of classical consonance definitions and emphasizing context-dependence.
Mazzola’s model allows a Kuratowski closure operator induced by a polarity, which defines an internal topology enabling algebraic-topological analysis of counterpoint structure. We close proving this construction generalizes to involutive morphisms. This categorical approach provides foundations for understanding both the historical evolution of contrapuntal practice and cross-cultural divergences in interval organization.
Key words and phrases:
topos theory, counterpoint, dichotomies, mathematical musicology
2020 Mathematics Subject Classification:
00A65, 18B05, 03G30
This work was partially supported by a grant from the Niels Hendrik Abel Board.
1. Introduction
Our intention in generalizing Mazzola’s counterpoint theory [2] via category and topos theory is to relativize the notion of consonance and dissonance, and to be able to apply contrapuntal techniques to other musical objects besides pitch. In particular, the topologization of the counterpoint model via a Kuratowski closure can be put in perspective within the general investigation of closure operators (see [4, 5], for example). More precisely, Mazzola does all the classical constructions of music in the functor category $R\mbox{-}\mathbf{Mod}_{fg},\mathbf{Set}]$ for some ring $R$. We think that a generalization by making $\mathbf{Set}$ any topos satisfying certain conditions and $R$-$\mathbf{Mod}_{fg}$ any (essentially) small category would make mathematical musicology gain more power of explanation. For example, the logic of the internal language of a topos might be more appropriate to reflect the development of counterpoint throughout history via the pseudocomplement. Our definitions are relative to a functor $F:\mathcal{M}\rightarrow\mathcal{E}$ where $\mathcal{M}$ is a (essentially) small category and $\mathcal{E}$ a suitable topos whose properties will be collected from section 2 to section 6 in order to justify why it should be a topos satisfying certain conditions.
The general plan of the article is to construct a categorical generalization of Mazzola’s original conception [10, Part VII], providing examples of certain gains we obtain from it, and thus refining our requirements on the ambient categories required to have a successful generalization; topoi appear as a satisfactory option. We close with a brief study of the Kuratowski operator introduced by Mazzola in order to “topologize” counterpoint in the generalized setting. We presuppose from the reader some familiarity with both counterpoint in general and Mazzola’s model in particular (see [2], [6], and [9] for general introductions on these topics).
We generalize the notion of polarity with the concept of quasipolarity in Section 2 in order to introduce the concept of dichotomy in Section 3. In Section 4 we introduce the concept of polarity proper. We present the notions of consonance and counterpoint symmetry in Section 5. All of these sections are accompanied by examples in order to illustrate the concepts. In Section 6, by their appearance in the definitions, we take stock of all the properties our suitable ambient category $\mathcal{E}$ should have. Finally, in Section 7 we categorize the notion of Kuratowski operator as introduced by Mazzola.
2. Quasipolarities
Let $\mathcal{E}$ be a suitable [^1] Below we will be more precise on what we need from it to be suitable. topos, $\mathcal{M}$ a small category, and $F:\mathcal{M}\longrightarrow\mathcal{E}$ a functor. Let $S$ be an object of $\mathcal{M}$. A quasipolarity for $F$ is a morphism $p:S\longrightarrow S$ of $\mathcal{M}$ satisfying the following conditions:
  1. i)
    The identity $p\circ p=\mathrm{id}_{S}$ holds.
  2. ii)
    The unique morphism from the initial object $\mathbf{0}$ of $\mathcal{E}$ to $F(S)$ is the equalizer (in $\mathcal{E}$) of the pair $F(p),\mathrm{id}_{F(S)}:F(S)\longrightarrow F(S)$.
Remark 2.1.
If the election of the functor $F$ is obvious (particulary if it is the forgetful functor), then we will omit it.
This definition was done thinking of the following example.
Example 2.2.
Take $\mathcal{E}=\textbf{Set}$, $\mathcal{M}=\mathbf{ModAf}_{\mathbb{Z}_{12}}$ (where $\mathbf{ModAf}_{R}$ denotes the category of modules over a commutative ring with affine transformations between them [^2] We prefer this notation instead of $\mathbf{Mod}_{R}$ used by Mazzola, since the latter is the standard notation for the category of $R$-modules and $R$-homomorphisms.), $F$ the forgetful functor from modules to sets, $S=\mathbb{Z}_{12}$, then $p=e^{2}.5$ is a quasipolarity.
Example 2.3.
The presence of the category $\mathcal{M}$ in the definition of quasipolarity is crucial. For instance, if $\mathcal{E}=\mathbf{Set}$ like in the previous example but now $\mathcal{M}=\mathbf{FinSet}$, $F$ is the inclusion functor and $S=\{0,1,\ldots,11\}$ (which coincides with $\mathbb{Z}_{12}$), then the two permutations
$p$ $=(0,2)(1,7)(3,5)(4,10)(6,8)(9,11),$
$q$ $=(0,1)(2,3)(4,5)(6,7)(8,9)(10,11)$
are quasipolarities; $p$ is exactly the same one as the one from the previous example as a function between sets, whereas $q$ does not come from an affine transformation.
3. Dichotomies
Let $p:S\longrightarrow S$ be a quasipolarity. A dichotomy relative to $p$ is a pair of monomorphisms
$\kappa:K\rightarrowtail F(S)\quad\text{and}\quad\delta:D\rightarrowtail F(S)$
in $\mathcal{E}$ such that
  1. i)
    The canonical morphism from the coproduct $K\amalg D$ (in $\mathcal{E}$) to $F(S)$ is an isomorphism.
  2. ii)
    The monomorphisms $F(p)\circ\kappa$ and $\delta$ represent the same subobject of $F(S)$.
We denote a dichotomy relative to $p$ with $(\kappa/\delta)_{p}$ or simply $(\kappa/\delta)$ if the election of $p$ is clear.
Example 3.1.
The subsets
$K=\{0,3,4,7,8,9\}\qquad\text{and}\qquad D=\{1,2,5,6,10,11\}$
of $\mathbb{Z}_{12}$ constitute a dichotomy $(K/D)_{e^{2}.5}$ for the polarity of the Example 2.2.
Proposition 3.2.
If $F:\mathcal{M}\rightarrow\mathbf{Set}$ is a functor with $\mathcal{M}$ a small category, then every quasipolarity $p:S\rightarrow S$ for $F$ has a dichotomy associated to it.
Proof.
Let $p:S\rightarrow S$ be a quasipolarity. Consider $Fp:FS\rightarrow FS$. If $FS$ is finite, it is easy to see that $p$ has a dichotomy associated to it.
Suppose $FS$ is infinite. Consider the following relation over $FS$: given $s,t\in FS$, we say $sRt$ if and only if $s=Fp(t)$. Let $R^{\prime}$ be the equivalence relation generated by $R$. Now, by the Axiom of Choice, form the $K$ by choosing one element from each equivalence class under the equivalence relation $R^{\prime}$. Let $D:=Fp(K)$. Then $Fp(D)=K$ and $FS=K+D$. ∎
It is known that the first and foremost consonances were the unison, the fifth and the fourth, while much later the “imperfect” thirds and sixth were added. In order to capitalize on the intuitionistic logic of a topos to model the historical development of counterpoint, we will broaden the notion of dichotomy as follows.
Definition 3.3.
A weak quasidichotomy relative to a quasipolarity $p:S\rightarrow S$ for a functor $F:\mathcal{M}\rightarrow\mathcal{E}$ is a $K\rightarrowtail FS$ satisfying $pK\subseteq K^{c}$, where $K^{c}$ is the pseudocomplement of $K$ with respect to $FS$. If $K$ is maximal then we say it is a quasidichotomy.
Example 3.4.
As reported by Sachs [16, p. 60], Johannes de Garlandia (among other 13th century theorists) considered as consonant (with varying degrees of “perfection”) what can be represented by the set $K=\{0,3,4,5,7\}$. This set does not correspond to a dichotomy, but it is a weak quasidichotomy, for there exist four involutive affine morphisms, namely $e^{1}.11$, $e^{6}.1$, $e^{9}.7$ and $e^{10}.5$, that map $K$ to a subset of the complement $K^{c}=\{1,2,6,8,9,10,11\}$.
The idea behind this definition is to construct the object $A$ determined by the formula
$\{u\subseteq FS:p(u)\subseteq u^{c}\}.$
This object would be a poset in $\mathcal{E}$ ordered by inclusion. If $\mathcal{E}$ is a topos satisfying some version of Zorn’s Lemma, we might have a quasipolarity relative to $p$ with something weaker than AC or IC. Thus the previous definition raises the following question: are there topoi satisfying some version of Zorn’s Lemma? The answer is yes; it is established in the following proposition.
Definition 3.5.
Let $\mathcal{E}$ be a topos and $A\in\mathcal{E}$. Given a term $w$ of power type $PA$ in the internal language of $\mathcal{E}$, define
$\operatorname{ub}(w):=\{x:\forall y\in w\,(y\leq x)\},$
i.e. $\operatorname{ub}(w)$ is the setlike term of upper bounds for $w$. Given a poset $(A,\leq)$ in $\mathcal{E}$, $A$ is said to be inductive in $\mathcal{E}$ if it satisfies
$\forall w\in PA.(w\text{ is a chain})\Rightarrow\exists x\in\operatorname{ub}(w).$
Proposition 3.6.
Assume Zorn’s Lemma holds in $\mathbf{Set}$, and let $\mathcal{E}$ be a localic $\mathbf{Set}$-topos. If $(A,\leq)$ is an inductive poset in $\mathcal{E}$, then there exists $a:1\rightarrow A$ which is (internally) a maximal element of $A$, i.e. satisfies $\forall x\in A.a\leq x\Rightarrow a=x$ ([7, D4.5.14]).
Given the notion of quasidichotomy, one could ask if in $\mathbf{Set}$ a quasidichotomy automatically becomes a dichotomy.
Proposition 3.7.
Let $F:\mathcal{M}\rightarrow\mathbf{Set}$ be a functor with $\mathcal{M}$ small. Let $p:S\rightarrow S$ be a quasipolarity for $F$. If $K\subseteq FS$ is a quasidichotomy relative to $p$, then $K$ is a dichotomy.
Proof.
Suppose $K$ is a quasidichotomy. Then $Fp(K)\subseteq K^{c}$ and $K$ is not contained in a larger set with the same property. Let us suppose $Fp(K)$ is not $K^{c}$. Then there exists $x\in K^{c}$ that is not in $Fp(K)$. Hence
$Fp(K)\subseteq K^{c}\setminus\{x\}=K^{c}\cap\{x\}^{c}=(K\cup\{x\})^{c}.$
Now, consider $K\cup\{x\}$. It cannot be the case that $Fp(K\cup\{x\})=Fp(K)\cup\{Fp(x)\}$ is contained in $(K\cup\{x\})^{c}$, since $K$ is maximal, thus $Fp(x)\in K\cup\{x\}$, but $Fp$ has no fixed points. Hence $Fp(x)\in K$. Therefore $F(pp)x=x$ is in $Fp(K)$, which is contradictory. Hence $Fp(K)=K^{c}$ for any maximal $K$. ∎
Remark 3.8.
Given the fact that when working in $\mathbf{Set}$, it is difficult (if not impossible) to distinguish between internal and external arguments, there are two possibilities for the proof of the previous proposition: it relies on the principle of excluded middle to arrive at the conclusion, i.e. it relies on the truth of $K^{c}\setminus Fp(K)=0$ or $\neg(K^{c}\setminus Fp(K)=0)$’ (equivalently, on the truth of $K^{c}\setminus Fp(K)=0$ follows from $\neg\neg(K^{c}\setminus Fp(K)=0)$’); or it relies on the strong witnessing of the internal logic of $\mathbf{Set}$ since $\mathbf{Set}$ is two-valued and supports split in it (see Propositions 4.31 and 4.32 in [3], and Axiom NS before 5.9). This says that a quasidichotomy relative to a quasipolarity for a functor whose codomain is a non-Boolean topos satisfying the version of Zorn’s Lemma in 3.6 may not be a dichotomy, or that a quasidichotomy relative to a quasipolarity for a functor whose codomain is a non-Boolean topos satisfying NS and the version of Zorn’s Lemma in 3.6 is a dichotomy.
We elaborate a little bit more on the second possibility. Suppose the functor $F$ has as codomain a topos $\mathcal{E}$ satisfying NS. Then for $K^{c}\setminus Fp(K)$ we would have $K^{c}\setminus Fp(K)=0$ or $K^{c}\setminus Fp(K)$ has a point, i.e. $K^{c}\setminus Fp(K)=0$ or $\neg(K^{c}\setminus Fp(K)=0)$; differently said using the turnstile of the deductive system of the internal logic of the topos:
(1) $\vdash_{\operatorname{Th}(\mathcal{E})}K^{c}\setminus Fp(K)=0\qquad\text{or}\qquad\vdash_{\operatorname{Th}(\mathcal{E})}\neg(K^{c}\setminus Fp(K)=0)$
because the system is complete (see [3, 4.32]). We suppose $\vdash_{\operatorname{Th}(\mathcal{E})}\neg(K^{c}\setminus Fp(K)=0)$ and we get a contradiction, but the system is consistent since the topos is two-valued. Hence we get
(2) $\not\vdash_{\operatorname{Th}(\mathcal{E})}\neg(K^{c}\setminus Fp(K)=0).$
So, from (1) and (2), we get $\vdash_{\operatorname{Th}(\mathcal{E})}K^{c}\setminus Fp(K)=0$. We didn’t use the principle of the excluded middle, neither outside nor inside the system, but NS. (Notice that $K^{c}\cap\{x\}^{c}=(K\cup\{x\})^{c}$ is valid in intuitionistic logic).
Example 3.9.
We give an example in a non-Boolean topos satisfying a version of Zorn’s Lemma as in 3.6 but not NS. Consider the Sierpiński topos $\mathbf{Set}^{\bullet\rightarrow\bullet}$, which is localic. Let $f:4\rightarrow 6$ be the inclusion function of $4=\{0,1,2,3\}$ into $6=\{0,\dots,5\}$. Now, let $p:4\rightarrow 4$ be the following composite of transpositions: $(0,2)(1,3)$, and let $q:6\rightarrow 6$ be that of $(0,2)(1,3)(4,5)$. The following diagram commutes:
[…]
Let $k=id_{2}$. Hence $\langle p,q\rangle(k)\subseteq k^{c}$, where $k^{c}$ is the inclusion of $\{2,3\}$ into $\{2,3,4,5\}$. It is clear that $k$ is a maximal subfunction of $f$ such that $\langle p,q\rangle(k)\subseteq k^{c}$, and that $\langle p,q\rangle(k)\neq k^{c}$.
4. Polarities
Let $(\kappa/\delta)$ be a dichotomy relative to a quasipolarity $p:S\longrightarrow S$. We say that the dichotomy is strong if $p$ is the unique quasipolarity such that $F(p)\circ\kappa$ and $\delta$ represent the same subobject; in that case, we say that $p$ is a polarity.
Example 4.1.
The dichotomy of the Example 3.1 is strong and its quasipolarity is a polarity.
Example 4.2.
The notion of polarity also stresses the importance of the choice of the category $\mathcal{M}$ and the functor [^3] A purely combinatorial relativization was explored in [1]. $F:\mathcal{M}\rightarrow\mathcal{E}$. For instance, if $\mathcal{M}_{1}=\mathbf{ModSAf}_{\mathbb{Z}_{12}}$, which is the subcategory of modules over a commutative ring with the morphisms restricted to those of the form $e^{u}.\pm 1$, taking
$(\kappa/\delta)=(\{0,2,3,4,7,8\}/\{1,5,6,9,10,11\}),$
we have that $p=e^{1}.-1$ is a polarity, but in the supercategory $\mathcal{M}_{2}=\mathbf{ModAf}_{\mathbb{Z}_{12}}$ it is not since
$e^{1}.-1(\kappa)=e^{9}.7(\kappa)=\delta.$
It is an open question whether every quasipolarity is a polarity.
5. Counterpoint symmetries
Definition 5.1 (Consonances).
Let $(\kappa/\delta)$ be a dichotomy. A consonance is a morphism $\xi$ (of $\mathcal{E}$) from the terminal object $\mathbf{1}$ of $\mathcal{E}$ to $K$, that is, a point of $K$. More generally, a generalized consonance can be defined as a morphism $\xi$ of $\mathcal{E}$ with codomain $K$.
Let $(\kappa/\delta)$ be a dichotomy with quasipolarity $p:S\longrightarrow S$. A counterpoint symmetry for a consonance $\xi:\mathbf{1}\longrightarrow K$ is an isomorphism $g:S\longrightarrow S$ of $\mathcal{M}$ such that
  1. i)
    the morphism $\kappa\circ\xi$ is a point of $F(g)\circ\delta$, or, more precisely, $\kappa\circ\xi$ factors through $F(g)\circ\delta$.
  2. ii)
    the identity $g\circ p=p\circ g$ holds, and
  3. iii)
    if an isomorphism $g^{\prime}:S\longrightarrow S$ of $\mathcal{M}$ satisfies i and ii, then there is a monomorphism from the meet [^4] We should keep in mind that a category can be seen as a generalization of the notion of partially ordered set; the meet of a pair of subobjects $m:S\rightarrowtail A$ and $m^{\prime}:S^{\prime}\rightarrowtail A$ can be obtained by means of the pullback of $m$ and $m^{\prime}$, whenever the category has pullbacks. $(F(g^{\prime})\circ\kappa)\wedge\kappa$ to the meet $(F(g)\circ\kappa)\wedge\kappa$ of $\mathcal{E}$.
Remark 5.2.
The subobjects $F(g)\circ\kappa$ and $F(g)\circ\delta$ of $F(S)$ represent the $g$-deformed consonances and dissonances, respectively.
Remark 5.3.
Since $\kappa\circ\xi$ is a deformed dissonance, then the tension induced by $g$ has to be relaxed towards a consonance that is also a $g$-deformed consonance. Thus the subobject $(F(g)\circ\kappa)\wedge\kappa$ corresponds to the admitted successors. More precisely, a consonance $\eta:\mathbf{1}\longrightarrow K$ is an admitted succesor of the consonance $\xi$ if the morphism $\kappa\circ\eta$ factorizes through $(F(g)\circ\kappa)\wedge\kappa$.
In the case when $\mathcal{E}=\mathbf{Set}$ and $\mathcal{M}=\mathbf{ModAf}_{\mathbb{Z}_{12}}$, the condition ii is exactly the same requirement stated by the classical model of Mazzola. The condition i captures the idea of revealing the tension between consonance and dissonances by making $\xi$ a $g$-deformed dissonance, while condition iii implies a maximum of artistic choices.
More generally, a counterpoint symmetry for a generalized consonance $\xi:E\longrightarrow K$ is an isomorphism $g:S\longrightarrow S$ of $\mathcal{M}$ satisfying that $\kappa\circ Im(\xi)$ factors through $F(g)\circ\delta$, plus the conditions ii and iii above. In this case, we require the existence of the image $Im(\xi)\rightarrowtail K$ of $\xi$.
Remark 5.4.
The notion of cantus firmus and discantus is obtained for general $R$-modules as follows. Let $R$ and $S$ be commutative rings. The restriction $r:R\to S$ gives rise to the functor
$S\otimes_{R}\underline{\ \ }:\mathbf{ModAf}_{R}$ $\to\mathbf{ModAf}_{S}$
$N$ $\mapsto S\otimes_{R}N,$
$f=e^{k}.f_{0}:N\to K$ $\mapsto S\otimes_{R}f=e^{1\otimes k}.S\otimes_{R}f_{0}.$
Recall now that the dual numbers for a ring $R$ is the quotient
$R[\epsilon]:=\frac{R[x]}{\langle x^{2}\rangle}=\{a+\epsilon.b:a,b\in R,\epsilon^{2}=0\}$
We have the restriction $i:R\to R[\epsilon]:a\mapsto a+\epsilon.0$, that allow us to construct, for an $R$-module $M$, the module $M[\epsilon]=R[\epsilon]\otimes_{R}M$, which is the dual numbers module with respect to $M$. As $R$-modules, we have
$M[\epsilon]\cong M\oplus M,$
and thus the first component of $(m_{1},m_{2})\in M[\epsilon]$ corresponds to the cantus firmus and $m_{2}$ is the interval that separates it from the discantus. The composition of this functor with the forgetful functor $F:\mathbf{ModAf}_{\mathbb{Z}_{12}[\epsilon]}\to\mathbf{Sets}$ recuperates the classical Mazzola model for counterpoint via counterpoint symmetries.
Example 5.5.
Let $S$ be finite set. We know that $2^{S}$ can be seen as a $\mathbb{Z}_{2}$-algebra (the so-called Boolean algebra) defining the product as intersection and addition as the symmetric difference
$A+_{2^{S}}B\equiv A\triangle B=(A\cup B)\setminus(A\cap B).$
In particular,
$p=e^{S}.1$
is a quasipolarity that coincides with the operation of calculating the complement in $S$, and it is the polarity of any family $\kappa$ of $2^{|S|-1}$ sets such that $T\in\kappa$ if, and only if,
$p(T)=S\triangle T=\complement T\notin\kappa.$
Thus we can construct the algebra $2^{S}[\epsilon]$ such that the cantus firmus and the interval are sets. The sets of consonances and dissonances in $2^{S}[\epsilon]$ are, of course,
$\kappa[\epsilon]=2^{S}+\epsilon.\kappa\quad\text{and}\quad\delta[\epsilon]=2^{S}+\epsilon.\delta.$
As in the case of the classical counterpoint model, it can be proved that the calculation of counterpoint symmetries can be reduced to cantus firmus 0 and the quasipolarity
$p_{0}=e^{\epsilon.S}.1.$
The affine morphism $g=e^{\epsilon.U}.(1+\epsilon.W)$ always commute with $p_{0}$, and for a consonance $\xi=\epsilon.K\in\kappa$ it will occur that it is a $g$-deformed dissonance if there is $C+\epsilon D\in\delta[\epsilon]$ such that
$\xi=\epsilon K=g(C+\epsilon.D)=C+\epsilon.(U\triangle D\triangle(W\cap C)),$
thus $C=0$ and
$K=U\triangle D.$
From this point on it is trivial to generalize Hichert’s algorithm [2, Algorithm 2.1] to complete the calculations.
With these preliminaries, we can obtain a rudimentary counterpoint theorem, i.e., the existence of admitted successors for any consonance.
Lemma 5.6.
Let $\mathcal{E}$ be a topos, and $K_{1},K_{2}$ be subobjects of $S\in\mathcal{E}$. If $K_{2}$ is a complemented subobject of $S$ and $K_{1}$ is not a subobject of $K_{2}$ then $K_{1}\cap K_{2}^{c}\neq 0$, where $K_{2}^{c}$ denotes the complement of $K_{2}$.
Proof.
If $K_{1}\cap K_{2}^{c}=0$, then, since $K_{2}$ is complemented, $K_{1}$ is a subobject of $K_{2}$, which is a contradiction. ∎
A sufficient condition for the existence of admitted successors is natural: that the effect of counterpoint symmetries is not to map all deformed consonances into consonances.
Proposition 5.7 (Counterpoint theorem).
Let $(K/D)$ be a dichotomy with quasipolarity $p:S\rightarrow S$, and $g$ be a counterpoint symmetry for some consonance. If $F(g)K$, the image of $K$ under $Fg$, is not a subobject of $D$, then $F(g)K\cap K\neq 0$.
Proof.
It follows by the previous lemma. ∎
Remark 5.8.
If the functor $F$ has $\mathbf{Set}$ as codomain and $S$ is finite, then if $F(g)K$ is different from $D$ (a weaker condition) we still have $F(g)K\cap K\neq\emptyset$.
The previous proposition gives us a condition to a possibility for admitted successors for consonances in a general ambient topos. However, this does not guarantee that $F(g)K\cap K$ has points. If we want that, we could consider the following property for a topos:
(NS) For every object $A$, either $A\cong 0$ or $A$ has a point. Equivalently, the topos is two-valued and supports split therein.
So this posits the question: are there topoi satisfying some version of Zorn’s Lemma and NS? The answer is yes:
Theorem 5.9.
Let $p:\mathcal{E}\rightarrow\mathbf{Set}$ be the canonical geometric morphism for a Grothendieck topos $\mathcal{E}$. Assume Zorn’s Lemma holds in $\mathbf{Set}$. If $(A,\leq)$ is an inductive poset in $\mathcal{E}$ and $p$ is precohesive over $\mathbf{Set}$, then there exists $a:1\rightarrow A$ which is (internally) a maximal element of $A$.
Proof.
Consider the external poset $C$ of chains in $A$, ordered by inclusion. As $C$ is an inductive poset in $\mathbf{Set}$ and $\mathcal{E}$ has arbitrary unions of subobjects since it is defined over $\mathbf{Set}$, $C$ has an external maximal element $A_{0}$. Hence
$A_{0}\in PA\wedge A_{0}\in C.$
Hence, since $A$ is inductive, we have $\exists x\in\operatorname{ub}(A_{0})$. That is, $\operatorname{ub}(A_{0})\rightarrow 1$ is epic. Hence $\operatorname{ub}(A_{0})\neq 0$. So, since $\mathcal{E}$ satisfies NS (see Motivation in [15]), $\operatorname{ub}(A_{0})$ has a point $a:1\rightarrow\operatorname{ub}(A_{0})$. It is clear that $A_{0}\cup\{a\}$ is a chain with $A_{0}\subseteq A_{0}\cup\{a\}$, but $A_{0}$ is maximal. Therefore
$A_{0}\cup\{a\}=A_{0}.$
This means $a\in A_{0}$. Whence $a$ is a maximal element of $A$. ∎
Example 5.10.
Some examples of non-Boolean presheaf topoi precohesive over $\mathbf{Set}$ are $\mathbf{Set}^{E}$, where $E$ is the monoid with exactly two elements: the unit and an idempotent element; the topos $\widehat{\Delta_{1}}$ of reflexive graphs; $(\mathbf{FinOrd}-0)^{op},\mathbf{Set}]$, where $\mathbf{FinOrd}$ is the category of finite ordinals with arrows all functions between them. More examples can be found in [12, 4.2 and 4.3] and in [13, 7.1 and 8.5]. Some examples of non-Boolean sheaf topoi precohesive over $\mathbf{Set}$ can be found in [13, 1.7 and 12.4].
Since we know that localic $\mathbf{Set}$-topoi satisfies some form of Zorn’s Lemma (see 3.6), we might want to know if there are localic $\mathbf{Set}$-topoi satisfying NS. This is not the case.
Proposition 5.11.
No localic $\mathbf{Set}$-topos satisfies NS.
Proof.
Let $\operatorname{Sh}(H)$ be a localic $\mathbf{Set}$-topos (i.e. $H$ is a locale). Every representable funtor in $\mathcal{O}(H)^{op},\mathbf{Set}]$ is a sheaf. But, by Yoneda, every representable $\mathcal{O}(H)(-,U)$ has no points except $\mathcal{O}(H)(-,1)$. ∎
6. Topology and closure operators
For this section we shall work with a category $\mathcal{E}$ with images and pullbacks [^5] This implies that, for each morphism of $\mathcal{E}$, the associated direct image functor is left adjoint to the associated inverse image functor, which in turn implies that the direct image functor preserves joins of subobjects (see [7, Lemma A.1.3.1])..
6.1. The operator induced by an involutive automorphism
The Kuratowski operator was introduced by Mazzola in counterpoint [2, Section 10.2.2] to topologize counterpoint intervals, so that homology could be calculated and related to the conditions determining the rules of first-species counterpoint. This construction can be generalized with no difficulty to the categorical case whenever we have an involutive morphism. Let $f:E\longrightarrow E$ be a morphism of $\mathcal{E}$ such that $f\circ f=\mathrm{id}_{E}$ [^6] Notice that, since $f$ is an isomorphism, $f$ preserves joins and meets without extra assumptions on $\mathcal{E}$; i.e., without $\mathcal{E}$ having images and pullbacks. However, we shall need those properties for the next subsection.. Define an operator on subobjects $M$ of $E$ by means of the equation
$\overline{M}:=M\vee f(M).$
We check this is indeed a Kuratowski closure.
  1. (1)
    If $m:M\rightarrowtail E$ is a subobject of $E$ in $\mathcal{E}$, then $M\leq M\vee f(M)=\overline{M}$.
  2. (2)
    If $m:M\rightarrowtail E$ is a subobject of $E$ in $\mathcal{E}$, then
    $\overline{\overline{M}}$ $=\overline{M}\vee f(\overline{M})$
    $=M\vee f(M)\vee f(M\vee f(M))$
    $=M\vee f(M)\vee f\circ f(M)$
    $=M\vee f(M)=\overline{M}.$
  3. (3)
    If $m:M\rightarrowtail E$ and $m^{\prime}:M^{\prime}\rightarrowtail E$ are subobjects of $E$ in $\mathcal{E}$, then
    $\overline{M\vee M^{\prime}}$ $=M\vee M^{\prime}\vee f(M\vee M^{\prime})$
    $=M\vee M^{\prime}\vee f(M)\vee f(M^{\prime})=\overline{M}\vee\overline{M^{\prime}}.$
6.2. The operator induced by an arbitrary endomorphism
As we have mentioned regarding the historical development of the notion of consonance and dissonance leading to Renaissance counterpoint (see [16, p. 60]), not always all intervals could be placed neatly into either category, but some approximative morphism is available. We analyse this case in this section.
Let $f:E\longrightarrow E$ be an arbitrary endomorphism of $\mathcal{E}$. Define an operator on subobjects $M$ of $E$ by means of the equation $\overline{M}:=M\vee f(M)$, where $f(M)$ denotes the image of $M$ under $f$. This operator satisfies the following properties:
  1. i)
    If $m:M\rightarrowtail E$ is a subobject of $E$ in $\mathcal{E}$, then
    $M\leq M\vee f(M)=\overline{M}.$
  2. ii)
    If $m:M\rightarrowtail E$ and $m^{\prime}:M^{\prime}\rightarrowtail E$ are subobjects of $E$ in $\mathcal{E}$, then
    $\overline{M\vee M^{\prime}}$ $=M\vee M^{\prime}\vee f(M\vee M^{\prime})$
    $=M\vee M^{\prime}\vee f(M)\vee f(M^{\prime})$
    $=\overline{M}\vee\overline{M^{\prime}}.$
In particular, this operator preserves the order on subobjects, though it needs not to be idempotent [^7] See [5, p. xiv] for another general discussion.. However, by iterating this operator, we can obtain an idempotent operator satisfying the previous properties, that is, a Kuratowski operator. In fact, if the lattice of subobjects of $E$ is complete (for example, if $\mathcal{E}$ is small-cocomplete), then the operator defined by the equation
$\overline{M}=\bigvee_{k\in\mathbb{N}}f^{k}(M)=M\vee f(M)\vee f^{2}(M)\vee\cdots$
satisfies the following properties:
  1. i)
    If $m:M\rightarrowtail E$ is a subobject of $E$ in $\mathcal{E}$, then $M\leq\overline{M}$.
  2. ii)
    If $m:M\rightarrowtail E$, then
    $\overline{\overline{M}}$ $=\bigvee_{k\in\mathbb{N}}f^{k}(\overline{M})$
    $=\bigvee_{k\in\mathbb{N}}f^{k}(\bigvee_{j\in\mathbb{N}}f^{j}(M))$
    $=\bigvee_{k\in\mathbb{N}}\bigvee_{j\in\mathbb{N}}f^{k+j}(M)$
    $=\bigvee_{k\in\mathbb{N}}f^{k}(M)=\overline{M}.$
  3. iii)
    If $m:M\rightarrowtail E$ and $m^{\prime}:M^{\prime}\rightarrowtail E$ are subobjects of $E$ in $\mathcal{E}$, then
    $\overline{M\vee M^{\prime}}$ $=\bigvee_{k\in\mathbb{N}}f^{k}(M\vee M^{\prime})$
    $=\bigvee_{k\in\mathbb{N}}f^{k}(M)\vee\bigvee_{k\in\mathbb{N}}f^{k}(M^{\prime})$
    $=\overline{M}\vee\overline{M^{\prime}}.$
In the case when $f^{n}=\mathrm{id}_{E}$ occurs for some $n\in\mathbb{N}$, this collapses to
$\overline{M}=M\vee f(M)\vee f^{2}(M)\vee\cdots\vee f^{n-1}(M).$
and of course, if $n=2$, then we obtain Mazzola’s closure operator from the previous subsection. Furthemore, given an endomorphism $p:S\longrightarrow S$ of $\mathcal{M}$, the definition above applies to the endomorphism $F(p):F(S)\longrightarrow F(S)$, and hence we have a Kuratowski operator on the subobjects of $F(S)$.
7. Some commentaries
The nature of the original definitions from Mazzola’s counterpoint model are set-theoretical. And rightly so, since Renaissance counterpoint consolidated the notion of consonance as a Boolean one: an interval is consonant or dissonant, never both. But in the framework of category theory, Heyting algebras are more natural, and could be used to reflect the historical development of counterpoint via the pseudocomplement. Another approach to non-Booleanness could be by means of fuzzy sets. However, if we take consonances to define a fuzzy set, that is, a function $\kappa:S\to[0,1]$, then the pseudocomplement
$\neg\kappa(x)=\begin{cases}0,&\kappa(x)>0,\\ 1,&\kappa(x)=0,\end{cases}$
defines, in particular, a crisp set, even if we apply it to a general fuzzy set. Further applications of this pseudocomplement yield only crisp sets, so we get back to Boolean complements (see [14, Theorem 2.1.9]). This is relevant since the progressive definition of the nature of consonance and dissonance that ended up with a stark separation between them with good contrapuntal properties is a well known musicological problem (see [16]); we must point out to [11] for a discussion of how this could be understood as a mathematical fact discovered by musical means. Hence, the intuitionistic logic of topoi can be used to take the first steps in order to solve this problem.
This led us to the issue of specifying a “suitable” topos for the purposes of musicology. We now recognize that we need a topos satisfying some form of Zorn’s Lemma if we want to guarantee the existence of quasidichotomies, and satisfying NS if we want to guarantee the existence of admitted successors for consonances.
A natural question for a future work is: if $S$ is a K-finite object of a topos $\mathcal{E}$, $p:S\rightarrow S$ an arrow in $\mathcal{E}$ such that $pp=1$ and the equalizer of $p$ and $1_{S}$ is 0, then is it true that for any subobject $A\rightarrowtail S$ maximal such that $pA$ is a subobject of $A^{c}$, we have $pA=A^{c}$?
Acknowledgment
We thank Guerino Mazzola for his valuable feedback on early versions of this paper.